OpenStudy (anonymous):

what is the fastest way to find the least common multiple of 32 and 56? (not using a calculator)

OpenStudy (anonymous):

if you know the gcd of the two numbers, it would be : \[\frac{(32)(56)}{\gcd(32,56)}\]

OpenStudy (anonymous):

gcd of 32 and 56 looks like 16, so: \[lcm(32,56) = \frac{(32)(56)}{16} = 2(56) = 112\]

OpenStudy (anonymous):

...which is wrong <.< ugh

OpenStudy (anonymous):

prime factorization then mutiply the greatest exponent

OpenStudy (anonymous):

ah, gcd(32,56) = 8, not 16, so it should really be: \[lcm(32,56) = \frac{(32)(56)}{8} = 4(56) = 224\] which is the right answer.